# Optimization under rare events: scaling laws for linear chance-constrained programs

### v1

@@ Abstract.
We consider a class of chance-constrained programs in which profit needs to be maximized while enforcing that a given adverse event remains rare. Using techniques from large deviations and extreme-value theory, we show how the optimal value scales as the prescribed bound on the violation probability becomes small and how convex programs emerge in the limit. We also use our results to show that the popular CVaR approximation is asymptotically optimal under light-tail assumptions, but it is sub-optimal in a heavy-tailed setting. Then, we apply point process techniques and random set theory to study the suboptimality gap in the Monte Carlo based scenario approach.
@@ 1. Introduction
Additional examples originate in the design of bridges and control of aircraft, cf. Shi and Holzapfel (2021), where the probability of a rare event is sometimes required to be below $\delta=10^{-9}$. For more examples, see Subramanyam (2020), Tong et al. (2022b).
@@
One method, introduced in Nemirovski and Shapiro (2006a), constructs a convex restriction of the feasible region by upper bounding the reliability constraint by the so-called conditional value at risk (abbreviated with CVaR; henceforth we refer to this as the CVaR approach). Another option, popular because of its ease of implementation, is the so-called scenario approach, introduced by Calafiore and Campi (2006). The scenario approach consists of replacing the probability constraint by several Monte Carlo-based scenarios, all of which need to satisfy the reliability constraint. If the undesirable event needs to be kept rare, the number of required samples is large, which can be prohibitive, though it is sometimes possible to reduce the number of samples using importance sampling techniques, see Nemirovski and Shapiro (2006b), Barrera et al. (2016), Blanchet et al. (2024).
Recently, several authors (see e.g. Subramanyam (2020), Nesti et al. (2019), Tong et al. (2022a)) have suggested a different method, namely to approximate the probability constraint using ideas from large-deviations theory. The approach can be described as follows. Suppose that ${\bf P}(G(x))\leq\delta$ is a safety constraint, $x$ is the optimization decision variable, and $G(x)$ is an undesirable event. A (heuristic) large-deviations approximation of ${\bf P}(G(x))$ can be obtained by replacing the integral of the density of an underlying probabilistic model (say $f$) over the set $G(x)$ by a supremum, i.e.:
${\bf P}(G(x))=\int_{G(x)}f(u)\mathrm{d}u\approx\sup_{u\in G(x)}f(u)=\exp\{-\inf_{u\in G(x)}\log f(u)\}.$ (1)
@@
The papers Subramanyam (2020), Nesti et al. (2019), and Tong et al. (2022a) focus on various computational aspects of this bi-level procedure. The analysis in those papers requires sufficient tractability for the underlying density $f$, and this is carried out in situations where (for example) the randomness is a mixture of Gaussian distributions or where the randomness arises from another specific light-tailed distribution. Further, this analysis focuses on developing approximations as in (1) that result in computationally tractable procedures.
Moving on to a more precise description of our contributions, this paper takes a different perspective: we derive rigorous approximations like (1) allowing both light and heavy-tailed distributions; in the latter case, the analogues of (1) can be harder to formulate. In turn, we consider a sequence of CC problems indexed by the safety parameter $\delta$, and complete the following technical tasks:
@@
    we identify conditions for which the latter approaches are asymptotically optimal or not.
@@
    A similar, but more nuanced conclusion, once again in the light-tailed environment, holds for the scenario approach: in this case, we show that the sequence of random linear programs, indexed by the number of scenarios, converges, after appropriate normalization, to a deterministic optimization problem as the number of scenarios is large. However, it must be emphasized that to match the behavior of ($\rm{CCP}_{\delta}$), the scenario approach requires a careful choice in the number of scenarios $k$ as function of $\delta$, which differs from the standard prescription provided by Calafiore and Campi (2006); see also (6) below. Again, feasibility for the original problem cannot be guaranteed in general. If we take the standard recommendation for $k$ as given in Calafiore and Campi (2006) to guarantee feasibility with high probability, the gap between the actual optimal profit and the optimal profit resulting from the scenario approach diverges. We refer to Theorem 3 and its discussion for details.
@@
    If the underlying random variables are heavy-tailed (more precisely, if their tails are of regular variation), we show (Theorem 4) that our basic chance-constrained problem converges to a convex optimization problem after appropriate renormalization. It turns out that (1) is not adequate here: essentially logarithmic asymptotics as in (1) are not enough; a more precise asymptotic estimate is needed. Fortunately, such an estimate is available in the heavy-tailed case, using multivariate regular variation.
@@
    A normalized sequence of CVaR constraints converges to a similar limit as the chance constraints, except with a different value in a certain constant. This makes CVaR asymptotically suboptimal in the heavy-tailed case. For example, if the underlying random variables have a power law with index $\alpha$, CVaR yields a value of the objective function which is a factor $1-1/\alpha$ of the actual optimum, as $\delta\downarrow 0$. A detailed statement is Theorem 5 below.
@@
    The scenario approach behaves even worse in the heavy-tailed case: the appropriately normalized sequence of random optimization problems converges to a limit, but this limit is still random, cf. Theorem 6 below. This is in contrast to methods such as sample-average approximation (also known) as empirical optimization in which, as the Monte Carlo sample size increases, the sample-based approximation converges to the desired solution. Our results suggest that in the heavy-tailed case the scenario approach must be modified to obtain asymptotically optimal solutions. We utilize concepts from point-process theory by Resnick (2007) and random set theory, see Molchanov (2017), to establish these results.
The rest of this paper is organized as follows. We define our model and state our assumptions in Section 2. In Section 3, we derive our results for light-tailed distributions. Our results for heavy-tailed distributions can be found in Section 4. In Section 5 we consider an example showing how it is possible to model dependence between random variables in our framework for light-tailed distributions, and how using approximations like (1) impacts feasibility. Proofs of our results can be found in Section 7, which also develops a general-purpose result that circumvents the use of epi-convergence.
@@ 2. Problem description
Basic formulation. Consider the problem
@@
where $x\in\mathbb{R}^{m}$ is an $m$-dimensional decision vector, $c$ is a vector with non-negative entries, $X$ is a compact convex subset of $[0,\infty)^{m}$ which contains the set $[0,h]^{m}$ for some $h>0$, and $L\geq 0$ is an $n$-dimensional random vector. The elements of $L$ are often referred to as risk factors and may be dependent. We assume there exist $n\times m$ matrices $A^{(i)},i=1,...,d$ with non-negative entries and set
$\phi(x,L):=\sup_{i=1,...,d}x^{T}A^{(i)}L.$ (2)
This formulation is general enough to model linear programs with random coefficients in the constraints.
As mentioned earlier, the parameter $\delta>0$ denotes the risk tolerance. We focus on the case where the reliability constraint is a rare event, i.e. $\delta\downarrow 0$. We aim to understand how the optimal profit scales as $\delta\downarrow 0$.
@@
[…]
$\limsup_{\delta\downarrow 0}{\bf P}(\phi(x_{\delta},L)>1)/\delta\leq 1.$ (3)
We call $(x_{\delta})_{\delta>0}$ asymptotically optimal if
@@
We now introduce two formulations that are popular alternatives to ($\rm{CCP}_{\delta}$); we are interested in whether their solutions are asymptotically optimal. The determination of asymptotic optimality is facilitated by Theorems 1 and 4 below, as they establish the asymptotic scaling of ($\rm{CCP}_{\delta}$) for light-tail and heavy-tail distributions, respectively.
@@
${\bf CVaR}(\phi(x,L))=\inf_{\tau\in{\mathbb{R}}}\Big{[}\tau+\frac{1}{\delta}{\bf E}[(\phi(x,L)-1-\tau)^{+}]\Big{]}.$ (5)
In particular, they propose to solve the optimization problem
@@
where $[k]:=\{1,\ldots,k\}$. To ensure feasibility of ($\rm{CCP}_{\delta}$) with high enough probability, $k$ needs to be chosen appropriately as $\delta\downarrow 0$. In Calafiore and Campi (2006) it is shown that the choice
$k(\delta)=\frac{2}{\delta}\log\frac{1}{\beta}+2d+\frac{2d}{\delta}\log\frac{2}{\delta}$ (6)
gives a solution that is feasible with probability at least $1-\beta$. This bound is valid for any function $\phi$. We show in Theorems 3 and 6 below that it leads to conservative solutions in our setting: letting $X^{k(\delta)}$ be a solution of ($\mathrm{SCP}_{k}$) with $k=k(\delta)$, then the random variable $c^{T}X^{k(\delta)}/v_{\delta}$ always converges to 0 in probability under our assumptions.
@@ 3. Light tails
In this section, we analyze ($\rm{CCP}_{\delta}$), ($\rm{CVaR}_{\delta}$), and ($\mathrm{SCP}_{k}$) for a large class of distributions guaranteeing that all moments of $L$ are finite. In particular, we assume that $Q(x):=-\log{\bf P}(L>x)$ is multivariate regularly varying: there exists a constant $\beta>0$, a slowly varying function $h$ and a non-negative continuous function $\lambda:[0,\infty)^{n}\mapsto\mathbb{R}$ such that for $x\in[0,\infty)^{n}$, with $q(r)=h(r)r^{\beta}$,
@@
The convergence as stated above is pointwise. However, it turns out to hold uniformly on compact sets because of regular variation, see Resnick (2007). Note that if we divide by $|x|$ and use that $h(r|x|)/h(r)\rightarrow 1$ as $r\rightarrow\infty$ for $|x|>0$, we obtain
@@
This family of distributions includes multi-variate normal distributions (for which $\beta=2$), exponential and gamma distributions ($\beta=1$) and Weibull distributions with a sub-exponential tail ($\beta<1$). The dependence structure can be rather arbitrary: the main form of dependence which is relevant is to which extent large values in the sample are correlated, which is encoded in $\lambda$. We give an example of $\lambda$ in Section 5.
We now apply large-deviations theory to approximate ($\rm{CCP}_{\delta}$) for small $\delta$. Set for $b\in[0,\infty)^{n}$
$I(b):=\inf_{x}\big{\{}\lambda(x):bx\geq 1\big{\}}=|b|^{-\beta}I(b/|b|).$ (9)
In Section 7.1 we show that the set $\{b:I(b)\geq 1\}$ is convex, compact, possesses a non-empty interior, and
$-\log{\bf P}(b^{T}L>r)\sim I(b)q(r).$ (10)
Here, $\sim$ means that the ratio of both functions converges to $1$ as $r\rightarrow\infty$. Next, we use the principle of the largest term (see e.g. Ganesh et al. (2004), Lemma 2.1) to approximate our chance constraint by
$-\log{\bf P}(\max_{i}y^{T}A^{(i)}L>r)\sim\min_{i}I(y^{T}A^{(i)})q(r)=:J(y)q(r),$ (11)
making the change of variable $x=y/r$. Now, consider the problem
maximize $\mbox{maximize }c^{T}y\hskip 28.45274pt{\mbox{ subject to }}y\geq 0\mbox{ and }J(y)\geq 1.$ ($\rm{LT}$)
Formulation ($\rm{LT}$) is motivated by making the change of variable $x=y/r$ in ($\rm{CCP}_{\delta}$), replacing the probability in ($\rm{CCP}_{\delta}$) by the approximation ${\bf P}(\max_{i}y^{T}A^{(i)}L>r)=\exp(-J(y)q(r)(1+o(1))$, and choosing $r=q^{-1}(\log 1/\delta)$. This ensures that the chance constraint, after taking logarithms, has the desired behavior. To make this work, we assume that $q$ is continuous for large $r$.
@@
[…]
To prove this result, a verification of the pointwise convergence property (11) is not sufficient. The standard way in the optimization literature to prove a result of the type of Theorem 1 is through epi-convergence, see Rockafellar and Wets (1998). In our proof, we circumvent this and follow a more direct approach; we refer to Section 7 for details.
@@
In contrast, ($\rm{CVaR}_{\delta}$) produces an asymptotically optimal sequence of solutions:
@@
$v_{{\bf CVaR},\delta}/v_{\delta}\rightarrow 1$ as $\delta\downarrow 0$.
The reason why this is true is that the overshoot $\phi(y^{*}/q^{-1}(\log 1/\delta),L)-1$, conditional on $\phi(y^{*}/q^{-1}(\log 1/\delta)/,L))>1$, converges to 00. In addition, the optimizing value of $\tau$ in (5) converges to 00 as $\delta\downarrow 0$. Together, these two properties imply that the constraint in the CVaR formulation ($\rm{CVaR}_{\delta}$) behaves asymptotically the same as the probability in ($\rm{CCP}_{\delta}$). This makes the CVaR formulation appealing when optimizing under rare event constraints in a light-tail setting.
To assess ($\mathrm{SCP}_{k}$) in the region $k\rightarrow\infty$, let $L^{(j)},j=1,...,k$ be $k$ i.i.d. samples of $L$. Using the insights from Theorem 1, this motivates us to consider a scaled problem. We make a similar change of variables $x=y/q^{-1}(\log k)$. We exploit that $\phi(y/r,L)=\phi(y,L)/r$ and consider
@@
[…]
To recover the optimal growth rate for the optimal value $v_{\delta}$ of ($\rm{CCP}_{\delta}$), we see from Theorem 1 that it is required that $k$ is proportional to $1/\delta$. As the choice $k(\delta)$ to guarantee feasibility diverges faster than $1/\delta$, the choice $k=1/\delta$ may lead to an asymptotically infeasible solution.
@@
The sampling approach leads to a random solution $Y^{(k)}$, which, after appropriate normalization, converges to the solution of ($\rm{LT}$). Apart from the computational challenges that play a role if $k$ is large, this procedure leads to a tension between feasibility and optimality: to get a feasible solution, the choice $k(\delta)$ leads to an optimal profit which is of smaller order than $v_{\delta}$ as $\delta\downarrow 0$.
Finally, the CVaR approach, while not as tractable as ($\rm{LT}$), and more conservative than ($\rm{CCP}_{\delta}$), is asymptotically optimal in a light-tail regime, reinforcing the fact that CVaR is one of the main approaches for solving chance constraint programs. In the next section, we will examine if these findings still hold if the underlying random variables are heavy-tailed.

### v2

@@ Abstract.
We consider a class of chance-constrained programs in which profit needs to be maximized while enforcing that a given adverse event remains rare. Using techniques from large deviations and extreme value theory, we show how the optimal value scales as the prescribed bound on the violation probability becomes small and how convex programs emerge in the limit. We use our results to analyze the performance of existing popular approaches in the rare-event regime. We show that the popular CVaR and sample approximations have optimality properties under light-tailed assumptions on the randomness, while they behave sub-optimal in a heavy-tailed setting. Our results are derived using large deviations theory, extreme value theory, process techniques, and random set theory.
@@ 1. Introduction
Additional examples originate in the design of bridges and control of aircraft, cf. Shi and Holzapfel (2021), where the probability of a rare event is sometimes required to be below $\delta=10^{-9}$. For more examples, see Subramanyam (2020), Tong et al. (2022).
@@
One method, proposed in Nemirovski and Shapiro (2006a), constructs a convex restriction of the feasible region by upper bounding the reliability constraint by the so-called conditional value at risk (abbreviated with CVaR; henceforth we refer to this as the CVaR approach, see also Rockafellar and Uryasev (2002)). Another option, popular because of its ease of implementation, is the so-called scenario approach, introduced by Calafiore and Campi (2006). The scenario approach consists of replacing the probability constraint by several Monte Carlo-based scenarios, all of which need to satisfy the reliability constraint. If the undesirable event needs to be kept rare, the number of required samples is large, which can be prohibitive, though it is sometimes possible to reduce the number of samples using importance sampling techniques, see Nemirovski and Shapiro (2006b), Barrera et al. (2016), Blanchet et al. (2024).
Recently, several authors (see e.g. Subramanyam (2020), Nesti et al. (2019), Tong et al. (2022)) have suggested a different method, namely to approximate the probability constraint using ideas from large-deviations theory. The approach can be described as follows. Suppose that ${\bf P}(G(x))\leq\delta$ is a safety constraint, $x$ is the optimization decision variable, and $G(x)$ is an undesirable event. A (heuristic) large-deviations approximation of ${\bf P}(G(x))$ can be obtained by replacing the integral of the density of an underlying probabilistic model (say $f$) over the set $G(x)$ by a supremum, i.e.:
${\bf P}(G(x))=\int_{G(x)}f(u)\mathrm{d}u\approx\sup_{u\in G(x)}f(u)=\exp\Big(-\inf_{u\in G(x)}\big\{\log 1/f(u)\big\}\Big).$ (1)
@@
The papers Subramanyam (2020), Nesti et al. (2019), and Tong et al. (2022) focus on various computational aspects of this bi-level procedure. The analysis in those papers requires sufficient tractability for the underlying density $f$, and this is carried out in situations where (for example) the randomness is a mixture of Gaussian distributions or where the randomness arises from another specific light-tailed distribution. Further, this analysis focuses on developing approximations as in (1) that result in computationally tractable procedures.
Moving on to a more precise description of our contributions, this paper takes a different perspective: we derive rigorous approximations like (1) allowing both light and heavy-tailed distributions. In the latter case, the analogues of (1) can be harder to formulate. In turn, we consider a sequence of CC problems indexed by the safety parameter $\delta$, and complete the following technical tasks:
@@
    we identify conditions for which the latter approaches are asymptotically feasible and optimal or not.
@@
    A similar, but more nuanced conclusion, once again in the light-tailed environment, holds for the scenario approach: in this case, we show that the sequence of random linear programs, indexed by the number of scenarios, converges after appropriate normalization to a deterministic optimization problem as the number of scenarios becomes large. To match the behavior of ($\rm{CCP}_{\delta}$), the scenario approach requires a careful choice in the number of scenarios $k$ as a function of $\delta$. The standard prescription provided by Calafiore and Campi (2006), see (6) below, allows for such a choice, providing a feasible and asymptotically optimal solution with probability tending to one.
@@
    If the underlying random variables are heavy-tailed (more precisely, if their tails are of regular variation), we show (Theorem 4) that our basic chance-constrained problem converges to a convex optimization problem after appropriate renormalization. The approximation (1) is not adequate here: essentially logarithmic asymptotics are not enough and a more precise asymptotic estimate is needed. We provide such an estimate using multivariate regular variation.
@@
    A normalized sequence of CVaR constraints converges to a similar limit as the chance constraints. However, CVaR yields a value of the objective function which is a constant factor away from the actual optimum as $\delta\downarrow 0$. This makes CVaR asymptotically suboptimal in the heavy-tailed case. A detailed statement is Theorem 5 below.
@@
    The scenario approach behaves poorly in the heavy-tailed case: the appropriately normalized sequence of random optimization problems converges to a limit, but this limit is still random and feasibility to the original problem cannot be guaranteed, cf. Theorem 6 below. This is in contrast to methods such as sample-average approximation (also known) as empirical optimization in which, as the Monte Carlo sample size increases, the sample-based approximation converges to the desired solution. Our results suggest that in the heavy-tailed case the scenario approach must be modified to obtain asymptotically both optimal and feasible solutions. We utilize concepts from point-process theory by Resnick (2007) and random set theory, see Molchanov (2017), to establish these results.
The rest of this paper is organized as follows. We define our model and state our assumptions in Section 2. In Section 3, we derive our results for light-tailed distributions. Our results for heavy-tailed distributions can be found in Section 4. In Section 5 we consider an example showing how it is possible to model dependence between random variables in our framework for light-tailed distributions, and how using approximations like (1) impacts feasibility. This example allows for explicit calculations. Afterwards, we verify our theoretical results with a numerical example on portfolio optimization in Section 6. We give a short reflection in Section 7, and present our proofs in Section 8, which also develops a general-purpose result that circumvents the use of epi-convergence.
@@ 2. Problem description
Notational conventions. Given vectors $x,y$ of identical dimension, the inequalities $x\geq y$ and $x>y$ are defined component wise, i.e. if $x>y$ if $x_{i}>y_{i}$ for all elements $i$. For an $m$-dimensional vector $x$ we denote $|x|=\sum_{i=1}^{m}|x_{i}|$ throughout the paper.
Basic formulation. Let $X$ be a compact convex subset of $0,\infty)^{m}$ for some $m\in\mathbb{N}$, and assume $X$ contains the set $0,h]^{m}$ for some $h>0$. Let $n\in\mathbb{N}$, and $\phi:X\times[0,\infty)^{n}\to\mathbb{R}$ be a function. Consider the problem
@@
where $L\geq 0$ is an $n$-dimensional random vector. The elements of $L$ are often referred to as risk factors and may be dependent. We assume $\phi$ can be written in terms of $m\times n$ matrices $A^{(i)},i=1,...,d$ with non-negative entries: we take
$\phi(x,L):=\max_{i=1,...,d}x^{T}A^{(i)}L.$ (2)
This formulation is general enough to model linear programs with random non-negative coefficients in the constraints. To see this, let $C$ be a random matrix with column vectors $C_{j}$, and $b$ be a random vector. We can write the matrix inequality $x^{T}C\leq b$ as $\sup_{i}x^{T}C_{i}/b_{i}\leq 1$. Next, define $L$ as the row vector consisting of all the $C_{i}/b_{i}$ and define the matrix $A^{(i)}$ consisting of only 0’s and $1$’s (in particular, the entries are nonnegative) in such a way that $A^{(i)}L=C_{i}/b_{i}$. We see also that the elements of $L$ are non-negative, if the elements of $C_{i}/b_{i}$ are nonnegative.
Our choice of $\phi$ implies that 0 is always a feasible solution, regardless of the value of $\delta$. In the context of linear random stochastic inequalities of the form $x^{T}AL\leq b$, if there exists an $x_{0}$ such that ${\bf P}(x_{0}^{T}AL\leq b)=1$, one can always make a change of variables by setting $z=x_{0}+x$ and considering $z^{T}AL\leq b-x_{0}^{T}AL$. In this case, $z=0$ is always feasible and our results would examine the speed of convergence of the objective function towards $c^{T}z_{0}$ as $\delta\downarrow 0$. In this note, we focus on the specific form (2) of $\phi$ to allow for short proofs. For a more general setting, we refer to the more recent paper by Deo and Murthy (2024).
As mentioned earlier, the parameter $\delta>0$ denotes the risk tolerance. We focus on the case where the reliability constraint is a rare event, i.e. $\delta$ is close to 0. We aim to understand how the optimal profit scales as $\delta\downarrow 0$.
@@
Let $v_{\delta}$ be the optimal value of $c^{T}x$ in ($\rm{CCP}_{\delta}$). Let $(x_{\delta})$ be a sequence of $m$-dimensional vectors such that $x_{\delta}\in X$ that is also asymptotically feasible, i.e.,
$\limsup_{\delta\downarrow 0}{\bf P}\big(\phi(x_{\delta},L)>1\big)/\delta\leq 1.$ (3)
We call $(x_{\delta})_{\delta>0}$ asymptotically optimal if it is asymptotically feasible and
@@
We now introduce two formulations that are popular alternatives to ($\rm{CCP}_{\delta}$); we are interested in whether their solutions are feasible and asymptotically optimal. The determination of asymptotic optimality is facilitated by Theorems 1 and 4 below, as they establish the asymptotic behavior of the optimal value $v_{\delta}$ of ($\rm{CCP}_{\delta}$) for light-tail and heavy-tail distributions, respectively.
@@
${\bf CVaR}(\phi(x,L))=\inf_{\tau\in{\mathbb{R}}}\Big\{\tau+\frac{1}{\delta}{\bf E}\big[(\phi(x,L)-\tau)^{+}\big]\Big\},$ (5)
where $x^{+}:=\max(0,x)$, see also Rockafellar and Uryasev (2002).
In particular, Nemirovski and Shapiro (2006a) propose to solve the optimization problem
@@
where $k]:=\{1,\ldots,k\}$. To ensure feasibility of ($\rm{CCP}_{\delta}$) with high enough probability, $k$ needs to be chosen appropriately as $\delta\downarrow 0$. In Calafiore and Campi (2006) it is shown that the choice
$k(\delta,\varepsilon)=\frac{2}{\delta}\log\frac{1}{\varepsilon}+2m+\frac{2m}{\delta}\log\frac{2}{\delta}$ (6)
gives a solution that is feasible with probability at least $1-\varepsilon$. This bound is valid for any function $\phi(x,L)$ that is convex in $x$. We show in Theorem 3 below that the choice $k=k(\delta,\varepsilon)$ matches the behavior of the optimal value in the light-tailed case if $\varepsilon$ does not tend to 0 with $\delta$ too quickly.
@@ 3. Light tails
In this section, we analyze ($\rm{CCP}_{\delta}$), ($\rm{CVaR}_{\delta}$), and ($\mathrm{SCP}_{k}$) for a large class of distributions guaranteeing that all moments of $L$ are finite. In particular, we assume that $Q(x):=-\log{\bf P}(L>x)$ is multivariate regularly varying: there exists a regularly varying function $q(r)=h(r)r^{\beta}$ for some index $\beta>0$ and slowly-varying function $h$ ($h(ar)/h(r)\rightarrow 1$ as $r\rightarrow\infty$ for any $a>0$), and a non-negative continuous function $\lambda:[0,\infty)^{m}\mapsto\mathbb{R}$ such that $\inf_{x\geq 0:|x|=1}\lambda(x)>0$, and
@@
The convergence as stated above is pointwise. However, it also holds uniformly on compact sets because of regular variation, see Resnick (2007). If we divide by $|x|$ and use that $h(r|x|)/h(r)\rightarrow 1$ as $r\rightarrow\infty$ for $|x|>0$, we obtain
@@
For technical simplicity, we assume that $q(r)$ is continuous and increasing so that its inverse is well-defined. We refer to Deo and Murthy (2024) for a more general setting.
This family of distributions includes multi-variate normal distributions ($\beta=2$), exponential and gamma distributions ($\beta=1$) and Weibull distributions with a sub-exponential tail ($\beta<1$). The dependence structure can be rather arbitrary: the main form of dependence which is relevant is to which extent large values in the sample are correlated, which is encoded in $\lambda$. We give an example in Section 5.
We apply large-deviations theory to analyze the optimal value $v_{\delta}$ of ($\rm{CCP}_{\delta}$) for small $\delta$. Set
$I(b):=\inf_{x\in[0,\infty)^{m}}\big\{\lambda(x):b^{T}x\geq 1\big\}=|b|^{-\beta}I(b/|b|),\qquad b\in[0,\infty)^{m},$ (9)
Here, $b^{T}$ will represent $y^{T}A^{(i)}$ for some $i\in\{1,\ldots,d\}$. We state a proposition on the role of $I$ that we prove in Section 8. We make the change of variable $x=y/r$, and write $f\sim g$ if $f(r)/g(r)\to 1$ as $r\rightarrow\infty$, or if $f(\delta)/g(\delta)\to 1$ as $\delta\downarrow 0$.
Proposition 1.
The set $\{b:I(b)\geq 1\}$ is convex, compact, and possesses a non-empty interior. The following asymptotics hold as $r\to\infty$,
$-\log{\bf P}\big(b^{T}L>r\big)$ $\sim I(b)q(r),$ (10)
$-\log{\bf P}\big(\max_{i}y^{T}A^{(i)}L>r\big)$ $\sim\min_{i}I\big(y^{T}A^{(i)}\big)q(r)=:J(y)q(r).$ (11)
Now, consider the following problem, closely related to ($\rm{CCP}_{\delta}$),
$\mbox{maximize }c^{T}y\hskip 28.45274pt{\mbox{ subject to }}y\geq 0\mbox{ and }J(y)\geq 1.$ ($\rm{LT}$)
Formulation ($\rm{LT}$) is motivated by making the change of variable $x=y/r$ in ($\rm{CCP}_{\delta}$) with $r=q^{-1}(\log 1/\delta)$, replacing the probability in ($\rm{CCP}_{\delta}$) by the approximation (11). This ensures that the chance constraint, after taking logarithms, has the desired behavior. The constraint $x\in X$ is no longer relevant in ($\rm{LT}$) as $r\rightarrow\infty$, since we assumed $0,h]^{m}\subseteq X$ for some $h>0$, and $x=y/r$.
@@
Let $y^{*}$ solve ($\rm{LT}$). Then $v_{\delta}\sim c^{T}y^{*}/q^{-1}(\log 1/\delta)$ as $\delta\downarrow 0$.
To prove this result, a verification of the pointwise convergence property (11) is not sufficient, as it does not imply convergence of the optima to the optimum of the pointwise limiting function. To circumvent this, we prove a general statement (Proposition 2 below) that we also invoke in the proofs of the subsequent theorems.
@@
In contrast, ($\rm{CVaR}_{\delta}$) produces an asymptotically optimal sequence of solutions. It is more conservative than ($\rm{CCP}_{\delta}$), and thus feasible. Moreover, it is optimal:
@@
$v_{{\bf CVaR},\delta}\sim v_{\delta}$ as $\delta\downarrow 0$.
We informally describe why Theorem 2 holds. Firstly, conditional on $\{\phi\big(y^{*}/q^{-1}(\log 1/\delta),L\big)>\tau\}$, the overshoot vanishes, i.e., $\phi\big(y^{*}/q^{-1}(\log 1/\delta),L\big)-\tau\to 0$ almost surely. Secondly, we show in Section 8.2 that we can approximate the optimum in ($\rm{CVaR}_{\delta}$) from below by setting $\tau$ in (5) close to 1. Together, these two properties imply that the constraint in the CVaR formulation ($\rm{CVaR}_{\delta}$) behaves asymptotically the same as the probability in ($\rm{CCP}_{\delta}$). This makes the CVaR formulation appealing when optimizing under rare event constraints in a light-tail setting.
To assess the scenario formulation ($\mathrm{SCP}_{k}$) in the region $k\rightarrow\infty$, let $\big(L^{(j)}:j=1,...,k)$ be $k$ i.i.d. samples of $L$. Motivated by the insights from Theorem 1, we consider a scaled problem. We make a similar change of variables $x=y/q^{-1}(\log k)$. We exploit that $\phi(y/r,L)=\phi(y,L)/r$ and consider
@@
Let $Y^{(k)}$ solve (${\rm{SCP}\textnormal{-}\rm{LT}}_{k}$). Then, as $k\to\infty$, $c^{T}Y^{(k)}/q^{-1}(\log k)\rightarrow c^{T}y^{*}$, almost surely, with $y^{*}$ a solution of ($\rm{LT}$). In particular, when $k=k(\delta,\varepsilon)$ defined in (6), and $\varepsilon=\varepsilon(\delta)$ is an arbitrary sequence such that $\delta\log(1/\varepsilon)\to 0$, then $c^{T}Y^{(k(\delta,\varepsilon))}/v_{\delta}\to 1$ almost surely.
The condition $\delta\log(1/\varepsilon)\to 0$ allows for some $\varepsilon=\varepsilon(\delta)\to 0$. For such choices, the sequence of solutions obtained from $k=k(\delta,\varepsilon)$ is feasible with probability tending to one, and asymptotically optimal.
@@
Both the sampling and the CVAR approaches provide solutions which scale like the optimal value $v_{\delta}$, making both approaches viable in the light-tailed case. The standard version of the sampling approach requires a large number of scenarios $k(\delta,\varepsilon)$. In recent work, it has been shown how to adapt the sampling approach and reduce $k$ by Choi et al. (2024).
The CVaR approach, while not as tractable as ($\rm{LT}$), and more conservative than ($\rm{CCP}_{\delta}$), is asymptotically optimal in a light-tail regime, reinforcing the fact that CVaR is one of the main approaches for solving chance constraint programs. In the next section, we will examine to what extent these findings hold if the underlying random variables are heavy-tailed.